« Project Status: 25 Feb 2007 (Sunday Morning) | Main | Project Status: 25 Feb 2007 (IOP7; Sunday Evening) »

Project Status: 25 Feb 2007 (IOP7; mid-day Sunday)

IOP7 is going strong. The freezing levels have remained relatively constant at ~3500 feet +/- (roughly, just above Alta (ATA) which is hovering a few tenths to 1 deg. above 0C). SMART-R crew reported an inch or so of wet snow overnight at Forest Hill (FHL) that has since melted; current precip at FHL is rain with occasional sleet. Rain continues to fall at lower elevations and Blue Canyon (BLU) has added 17" of snow since 0Z. SMART-R shows a band of precip (w/45 dBZ echoes embedded) heading towards the ARB, out at about 50 km range (also visible on KDAX).

The latest forecast calls for freezing levels to remain in the 3300'-3700' range until FROPA which is expected at ~0Z Feb 27th (~4pm Monday, local time) at which point they will drop. The IOP7 end game is a question mark at this time as the models disagree about what will transpire from Tuesday night into Wednesday. No change in the operational plan for IOP7 is required at this time, we'll revisit this issue on Monday's daily status call at 11:30 PST.

Instrument Status Issues:
SMART-R crew reports that the radar has been running well - no issues at this time. Overnight snow was light and wet and did not stick to the dish. Soundings in Reno are going according to plan. There were a few minor issues with the Sloughhouse (SHS) soundings, but are generally OK. No one is sure what is happening with the Oakland soundings (supplemental or otherwise), we'll look into it. The SHS profiler is still acting up and early this afternoon, a technician will attempt to diagnose the problems remotely.

T. Schneider

TrackBack

TrackBack URL for this entry:
/cgi-bin/mt/mt-tb.cgi/280